If a current of 0.5 amperes flows through a metallic wire for two hours, then how many electrons flow through the wire?

Grade:upto college level

1 Answers

Aarti Gupta
9 years ago
Current,I = 0.5A
Time ,t = 2 hr
n , no. of e- =
Sol.
we know that
I = Q/t
where ,Q = charge
Q = I *t
= 0.5* 2
Q = 1C
NOW,
Q = ne
n = Q/e
= 1/1.6 *10-19
n = 0.625 * 1019
or,
n = 6.25 * 1018 electrons